# Improving Plasticity in Online Continual Learning via Collaborative Learning

### v1

@@ 4.2 Collaborative Continual Learning
where $(X,y)$ is the data-label pair, $\mathcal{L}_{cls}(\cdot)$ is the classification loss in the baseline method CCL adapts to, $D_{KL}(\cdot)$ is the Kullback-Leibler divergence, $\lambda_{1}$ and $\lambda_{2}$ are balancing hyperparameters and $\tau$ is the temperature hyperparameter. Note that the network $\theta^{2}$ should be trained with $\mathcal{L}_{CCL}^{2}$, respectively.
@@ 4.3 Distillation Chain
The main motivation of DC originates from the idea of entropy regularization-based optimization strategies, like label smoothing [44], knowledge distillation [49], and confidence penalty [36], where we find that overconfidence will hurt performance in non-continual scenarios. As shown in the supplementary material, we observed a similar phenomenon in online CL. To tackle the problem, DC uses data augmentation strategies to generate samples with different levels of difficulty and produces logit distribution with different confidence. The distillation from less confident predictions to more confident predictions weakens the overall confidence of the network and benefits the performance by improving the generalization capability.
@@
where $X_{i}$ is the augmentation of input sample $X$ after $i$ augmentation steps. More discussion about why and how DC works can be found in the supplementary material.
4.4 Apply CCL-DC to online CL methods
# model1: student model
# model2: teacher model
# optim1: optimizer for student model
# cls: classification loss in baseline
for x, y in dataloader:
# Baseline loss
loss_baseline = criterion_baseline(model1, x, y)
[…]
x1 = geometric_distortion(x)
x2 = RandAugment(x1, N, M)
x3 = RandAugment(x2, N, M)
[…]
ls, ls1, ls2, ls3 = model1(x, x1, x2, x3)
lt, lt1, lt2, lt3 = model2(x, x1, x2, x3) # no grad
[…]
loss_ccl = kl_div(ls/t, lt/t) # temperature t
loss_dc = kl_div(ls/t, lt1/t) + kl_div(ls1/t, lt2/t) + kl_div(ls2/t, lt3/t)
[…]
loss = loss_baseline + loss_ours
[…]
loss.backward()
optim1.step()
@@ 5 Experiments
Table 1: Average Accuracy (%, higher is better) on four benchmark datasets with difference memory buffer size $M$, with and without our proposed CCL scheme. The result of our method is given by the ensemble of two peer models. All values are averages of 10 runs.
@@
Table 2: Learning Accuracy (%, higher is better) on four benchmark datasets with difference memory buffer size $M$, with and without our proposed CCL scheme. The result of our method is given by the ensemble of two peer models. All values are averages of 10 runs.
5.1 Experimental Setup
Datasets.
We use four image classification benchmark datasets to evaluate the effectiveness of our method, including CIFAR-10 [27], CIFAR-100 [27], TinyImageNet [28], and ImageNet-100 [24]. More detailed information about the dataset split and task allocation is given in the supplementary material.
Baselines.
Implementation details.
5.2 Results and Analysis
Final average accuracy.
Table 1 presents the results of average accuracy (AA) at the end of the training on four datasets. As indicated in Sec. 4, to fully take advantage of collaborative learning, we show the results with the ensemble of two models, with the independent model performance available in Sec. 6. Generally, the ensemble method provides about 1% additional accuracy compared to independent inference. For all datasets, memory size $M$, and baseline methods, applying CCL-DC constantly improves the performance by a large margin. Notably, even for state-of-the-art methods like GSA and OnPro, we can still gain significant performance when incorporating CCL-DC.
@@ 5.2 Results and Analysis
Table 3: Relative Forgetting (%, lower is better) on four benchmark datasets with difference memory buffer size $M$, with and without our proposed CCL scheme. The result of our method is given by the ensemble of two peer models. All values are averages of 10 runs.
More interestingly, for almost all settings with different memory buffer sizes $M$, the performance gain tends to be a constant on a relative basis. For example, CCL-DC can boost the performance of ER on Tiny-ImageNet from 10.82 to 16.56 when $M$=2k, which is a 53.0% performance gain on a relative basis. The performance gain is 53.4% and 52.7% when $M$=5k and $M$=10k respectively. This indicates that we can achieve a decent performance gain regardless of the memory buffer size, and it shows the scalability of our method to different resource conditions.
@@
As mentioned in Sec. 3, we evaluate the plasticity and stability of different continual learners with Learning Accuracy and Relative Forgetting, respectively. Table 2 shows the plasticity metric on four datasets. For all settings, CCL-DC constantly improves the model plasticity by a large margin. For model stability, as indicated by RF in Table 3, models trained with CCL-DC are comparable with the baselines under most cases. ER-ACE is an exception as its plasticity is unexpectedly low, especially on TinyImagenet. Also, the stability of ER-ACE is compromised when incorporating CCL-DC. We will explain the reason for this unexpected phenomenon in the supplementary material.
@@ 5.3 Ablation Studies
We also evaluate the effectiveness of DC’s strategy of distilling from harder samples to easier samples in collaborative learning manner. As shown in Table 5, we compared it with other distillation strategies. The result shows that the distillation scheme of DC constantly outperforms other schemes. Extra experiments explaining the working mechanism of DC can be found in the supplementary material.
@@ 6 Discussions
Improving feature discrimination.
@@
Another advantage of CCL-DC is its ability to enhance the feature discrimination of continual learners. Fig. 4 illustrates the t-SNE visualization [46] of the memory data’s embedding space at the end of the training. We can see that the feature representation of the method with CCL-DC is more discriminative compared with the baseline.
@@
Moreover, we can evaluate the feature discrimination using the clustering methods. Following [31], we remove the final FC classifier and use Nearest-Class-Mean (NCM) [31] classifier with intermediate representations. Table 6 demonstrates that CCL-DC can greatly enhance the NCM accuracy, which evidences the capability of CCL-DC in improving feature discrimination.
Alleviating shortcut learning.
Shortcut learning [18] is another commonly observed issue that hinders the generalization capability of continual learners [48]. In Fig. 5, we use GradCAM++ [8] on the training set of ImageNet-100 (M=5k) at the end of the training of ER and GSA. Although both ER and GSA make correct predictions, we observed that they focus on irrelevant objects, which indicates a tendency toward shortcut learning. Also, we can see that by integrating CCL-DC, the shortcut learning can be greatly alleviated.
Method Ind. Acc. ↑ Ens. Acc. ↑
ER + Ours 43.58±1.05 44.45±1.04
DER++ + Ours 42.79±1.38 43.52±1.35
ER-ACE + Ours 44.15±1.05 45.14±1.00
OCM + Ours 42.39±1.36 43.34±1.51
GSA + Ours 43.84±1.34 44.74±1.32
OnPro + Ours 41.18±0.83 41.89±0.82
Table 7: Comparison of the final average accuracy achieved through independent inference and the use of the ensemble method on CIFAR-100 (M=2k). Independent accuracy (Ind. Acc.) is calculated by averaging the accuracy of two networks in CCL-DC. All values are averaged over 10 runs.
@@
Independent network performance.
Although the ensemble method gives extra performance at inference time, by averaging the logit output of two networks in CCL-DC, it also doubles the computation. In some cases, computational efficiency becomes more crucial during inference. Continual learners trained with CCL-DC are also able to do inference independently, albeit with a slight performance drop compared with ensemble inference. Table 7 illustrates the accuracy achieved through independent inference. It is evident that the performance loss in independent inference, when compared to ensemble inference, is minimal (approximately 1%).
7 Conclusion
In this paper, we highlight the significance of plasticity in online CL, which has been neglected in prior research. We also establish the quantitative link between plasticity, stability, and final accuracy. The quantitative relationship sheds light on the future direction of online CL research. Based on this, we introduce collaborative learning into online CL and propose CCL-DC, a strategy that can be seamlessly integrated into existing online CL methods. Extensive experiments show the effectiveness of CCL-DC in boosting plasticity and subsequently improving the final performance.
Supplementary Material
8 Working mechanism of DC
Overconfidence will hurt performance.
In conventional supervised learning, it is well established that excessive confidence can harm the generalization ability of deep models. To tackle this issue, many methods like label smoothing [44], knowledge distillation [49], and confidence penalty [36] are proposed.
In online CL, we find a counter-intuitive, yet similar phenomenon. As shown in Fig. 6, we manually stagnate a peer network $\theta^{2}$ in CCL (i.e., Kaiming initialize and stay untrained), and train the network $\theta^{1}$ with both baseline loss and the distillation loss from untrained network $\theta^{2}$, as illustrated in Eq. 9. And the experimental result in Table 8 shows a decent performance gain compared with independent training. The distillation from the prediction of untrained network $\theta^{2}$ to network $\theta^{1}$ serves as a regularization to suppress the overall confidence, and the performance gain under this condition makes it evident that, in continual scenarios, overconfidence will also harm the performance.
Figure 6: Conceptual diagram of the CCL framework, when distill from an untrained network $\theta^{2}$ to suppress the confidence of network $\theta^{1}$.
Dataset CIFAR-100
Memory Size $M$ M=1000 M=2000 M=5000
ER 24.47±0.72 31.89±1.45 39.41±1.81
ER + CCL($\theta^{2}$ untrained) 27.07±1.20 34.84±0.64 41.15±1.16
Table 8: The performance of network $\theta^{1}$ when distill from untrained network $\theta^{2}$ on CIFAR-100 (M=2k). All numbers are average over 10 runs.
Learned entropy regularization with DC.
As indicated in Sec. 4, the main idea of DC is the distillation from less confident predictions to more confident predictions. Similar to other entropy regularization strategies like label smoothing, knowledge distillation, and confidence penalty in non-continual scenarios, this distillation weakens the overall confidence of the network and benefits the performance by improving the generalization capability. To show the effect of the regularization by DC, we compared the confidence of models trained with and without CCL-DC. We measure the confidence with the entropy of predictions. Fig. 7 shows the entropy of prediction produced by the ER (baseline) with and without CCL-DC on CIFAR-100 (M=2k). The entropy is calculated at the end of training and averaged across all samples in the training set. Moreover, we also calculate the entropy of models trained with CCL-DC, by forwarding images of different difficulties (from $X_{0}$ to $X_{3}$ in Eq. 10) in the augmentation chain. The experimental results show the effect of entropy regularization of DC. Also, we believe that the working mechanism of DC has a connection to entropy regularization based methods like label smoothing.
9 Extra Experiments
Impact of the number of augmentation stages in DC.
T-SNE visualization on other baselines.
In Sec. 6, we show the t-SNE visualization on memory data at the end of training for ER and ER + Ours, and we prove the advantage of CCL-DC in improving feature discrimination. In Fig. 11, we show the t-SNE results for the remaining baselines. From the figure, we can see that even if the baseline is well-trained with state-of-the-art methods like OCM, GSA, and OnPro, CCL-DC can still significantly improve feature discrimination.
Classification loss curve on other baselines.
@@ 11 Experiment Details
All the experiments in our work are conducted on NVIDIA A100 GPUs. Fig. 13 shows the training time for each method with and without CCL-DC on CIFAR-100 (M=5k).

### v2

@@ 4.2 Collaborative Continual Learning
where $(X,y)$ is the data-label pair, $\mathcal{L}_{cls}(\cdot)$ is the classification loss in the baseline method CCL adapts to, $D_{KL}(\cdot)$ is the Kullback-Leibler divergence, $\lambda_{1}$ and $\lambda_{2}$ are the balancing hyperparameters and $\tau$ is the temperature hyperparameter. Note that the network $\theta^{2}$ should be trained with $\mathcal{L}_{CCL}^{2}$, respectively.
@@ 4.3 Distillation Chain
As shown in Table 2, we observed that overconfidence will hurt the performance in continual training. To tackle the problem, DC uses data augmentation strategies to generate samples with different levels of difficulty and produces logit distribution with different confidence. The distillation from less confident predictions to more confident predictions weakens the overall confidence of the network and benefits the performance by improving the generalization capability.
@@
where $X_{i}$ is the augmentation of input sample $X$ after $i$ augmentation steps.
{minted}
python # model1: student model # model2: teacher model # optim1: optimizer for student model # cls: classification loss in baseline for x, y in dataloader: # Baseline loss loss_baseline = criterion_baseline(model1, x, y)
# DC Augmentation x1 = geometric_distortion(x) x2 = RandAugment(x1, N, M) x3 = RandAugment(x2, N, M)
# CCL-DC loss ls, ls1, ls2, ls3 = model1(x, x1, x2, x3) lt, lt1, lt2, lt3 = model2(x, x1, x2, x3) # no grad
loss_cls = cls(ls, y) + cls(ls1, y) + cls(ls2, y) + cls(ls3, y) loss_ccl = kl_div(ls/t, lt/t) # temperature t loss_dc = kl_div(ls/t, lt1/t) + kl_div(ls1/t, lt2/t) + kl_div(ls2/t, lt3/t)
loss_ours = lam1*loss_cls + lam2*(loss_ccl + loss_dc) loss = loss_baseline + loss_ours
optim1.zero_grad() loss.backward() optim1.step()
@@
Table 3: Average Accuracy (%, higher is better) on four benchmark datasets with difference memory buffer size $M$, with and without our proposed CCL-DC scheme. The result of our method is given by the ensemble of two peer models. All values are averages of 10 runs.
@@
Table 4: Learning Accuracy (%, higher is better) on four benchmark datasets with difference memory buffer size $M$, with and without our proposed CCL-DC scheme. The result of our method is given by the ensemble of two peer models. All values are averages of 10 runs.
4.4 Apply CCL-DC to online CL methods
5 Experiments
@@
Table 5: Relative Forgetting (%, lower is better) on four benchmark datasets with difference memory buffer size $M$, with and without our proposed CCL-DC scheme. The result of our method is given by the ensemble of two peer models. All values are averages of 10 runs.
5.1 Experimental Setup
Datasets.
We use four image classification benchmark datasets to evaluate the effectiveness of our method, including CIFAR-10 [25], CIFAR-100 [25], TinyImageNet [26], and ImageNet-100 [22, 15, 11]. More detailed information about the dataset split and task allocation is given in the supplementary material.
Baselines.
Implementation details.
5.2 Results and Analysis
Final average accuracy.
Table 3 presents average accuracy (AA) at the end of the training on four datasets. As indicated in Sec. 4, to fully take advantage of collaborative learning, we show the results with the ensemble of two models, with the independent model performance available in the supplementary material. Generally, the ensemble method provides about 1% additional accuracy compared to independent inference. For all datasets, memory size $M$, and baseline methods, applying CCL-DC constantly improves the performance by a large margin. Notably, even for state-of-the-art methods such as GSA and OnPro, we can still gain significant performance when incorporating CCL-DC.
More interestingly, for almost all settings with different memory buffer sizes $M$, the performance gain tends to be a constant on a relative basis. For example, CCL-DC can boost the performance of ER on Tiny-ImageNet from 10.82 to 16.56 when $M$=2k, which is a 53.0% performance gain on a relative basis. The performance gain is 53.4% and 52.7% when $M$=5k and $M$=10k, respectively. This indicates that we can achieve a decent performance gain regardless of the memory buffer size, and it shows the scalability of our method to different resource conditions.
@@ 5.2 Results and Analysis
As mentioned in Sec. 3, we evaluate the plasticity and stability of different continual learners with LA and RF, respectively. Table 4 shows the plasticity metric on four datasets. For all settings, CCL-DC constantly improves the model plasticity by a large margin. For model stability, as indicated by RF in Table 5, models trained with CCL-DC are comparable with the baselines under most cases. ER-ACE is an exception as its plasticity is unexpectedly low, especially on TinyImagenet. Also, the stability of ER-ACE is compromised when incorporating CCL-DC. We will explain the reason for this unexpected phenomenon in the supplementary material.
@@ 5.3 Ablation Studies
We also evaluate the effectiveness of DC’s strategy of distilling from harder samples to easier samples in collaborative learning manner. As shown in Table 7, we compared it with other distillation strategies. The result shows that the distillation scheme of DC constantly outperforms other schemes.
Self-Distillation Chain in an independent model.
While DC was originally designed to incorporate CCL, it is possible to conduct DC’s strategy within an independent model. We name such a strategy as the Self-Distillation Chain (SDC). Similar to different schemes of DC, SDC can be implemented in two ways: distillation from easier samples to harder samples, and distillation from harder samples to easier samples. As shown in Table 8, both strategy gives extra final performance and plasticity, while the latter strategy benefits the performance more. Moreover, incorporating DC with CCL (i.e., ours) further improves the accuracy.
@@
Method Distillation scheme Acc. ↑ LA ↑
ER + SDC Easy to hard 35.00±1.31 56.67±0.84
ER + SDC Hard to easy 41.31±1.25 68.62±0.60
ER + CCL-DC Easy to hard 40.95±0.97 60.03±0.98
ER + CCL-DC Hard to easy (Ours) 44.45±1.04 70.86±0.72
Table 8: Final average accuracy when conducting multiview distillation within the same model on CIFAR-100 (M=2k). All values are averaged over 10 runs.
@@ 6 Discussions
Alleviating overconfidence.
To show the effect of DC in alleviating overconfidence, we compared the confidence of models trained with and without CCL-DC. We measure the confidence with the entropy of predictions. Fig. 5 shows the entropy of prediction produced by the ER (baseline) with and without CCL-DC on CIFAR-100 (M=2k). The entropy is calculated at the end of training and averaged across all samples in the training set. Moreover, we also calculate the entropy of models trained with CCL-DC, by forwarding images of different difficulties (from $X_{0}$ to $X_{3}$ in Eq. 10) in the augmentation chain. The experimental results show the effect of DC in suppressing overall confidence.
7 Conclusion
In this paper, we highlighted the significance of plasticity in online CL, which has been overlooked in prior research when compared to stability. We also established the quantitative link between plasticity, stability, and final accuracy. The quantitative relationship shed light on the future direction of online CL research. Based on this, we introduced collaborative learning into online CL and proposed CCL-DC, a strategy that can be seamlessly integrated into existing online CL methods. Extensive experiments showed the effectiveness of CCL-DC in boosting plasticity and subsequently improving the final performance.
Supplementary Material
8 Extra Experiments
Impact of the number of augmentation stages in DC.
T-SNE visualization.
@@ 8 Extra Experiments
Another advantage of CCL-DC is its ability to enhance the feature discrimination of continual learners. Fig. 10 illustrates the t-SNE visualization of the memory data’s embedding space at the end of the training. We can see that the feature representation of the method with CCL-DC is more discriminative compared with the baseline.
Classification loss curve on other baselines.
Independent network performance.
Although the ensemble method gives extra performance at inference time, by averaging the logit output of two networks in CCL-DC, it also doubles the computation. In some cases, computational efficiency becomes more crucial during inference. Continual learners trained with CCL-DC are also able to do inference independently, albeit with a slight performance drop compared with ensemble inference. Table 9 shows the accuracy achieved through independent inference. It is evident that the performance loss in independent inference, when compared to ensemble inference, is minimal (approximately 1%).
Dataset CIFAR10 CIFAR100 Tiny-ImageNet ImageNet-100
Memory Size $M$ 500 1000 1000 2000 5000 2000 5000 10000 5000
ER + Ours (Ind.) 65.66±2.35 73.37±1.70 32.97±1.06 43.58±1.05 52.96±1.16 16.32±1.58 28.68±1.20 37.14±0.93 41.82±1.54
ER + Ours (Ens.) 66.43±2.48 74.10±1.71 33.43±1.06 44.45±1.04 53.81±1.16 16.56±1.63 29.39±1.23 37.73±0.85 43.11±1.49
DER++ + Ours (Ind.) 68.15±1.40 73.56±1.12 33.81±0.90 42.79±1.38 52.04±0.81 11.11±1.53 21.47±1.93 27.37±2.64 44.22±1.25
DER++ + Ours (Ens.) 68.79±1.42 74.25±1.10 34.36±0.89 43.52±1.35 52.95±0.86 10.99±1.39 21.68±1.94 28.01±2.46 45.70±1.32
ER-ACE + Ours (Ind.) 69.35±1.24 74.86±1.06 36.34±1.08 44.15±1.05 52.94±0.44 17.99±1.56 25.69±2.00 31.69±1.69 43.92±1.71
ER-ACE + Ours (Ens.) 70.08±1.38 75.56±1.14 37.20±1.15 45.14±1.00 53.92±0.48 18.32±1.49 26.22±2.01 32.23±1.70 45.15±1.94
OCM + Ours (Ind.) 73.00±0.88 76.66±1.38 34.02±1.22 42.39±1.36 50.19±1.36 22.53±1.28 32.16±0.96 38.02±0.94 41.71±1.07
OCM + Ours (Ens.) 74.14±0.85 77.66±1.46 35.00±1.15 43.34±1.51 51.43±1.37 23.36±1.18 33.17±0.97 39.25±0.88 43.19±0.98
GSA + Ours (Ind.) 68.10±1.58 74.78±1.27 35.14±1.40 43.84±1.34 54.29±1.10 16.53±1.62 27.57±1.61 36.12±1.59 43.27±1.05
GSA + Ours (Ens.) 68.91±1.68 75.78±1.16 35.56±1.39 44.74±1.32 55.39±1.09 16.70±1.66 28.11±1.70 37.13±1.75 44.28±1.16
OnPro + Ours (Ind.) 73.65±2.16 77.84±1.33 34.20±1.12 41.18±0.83 49.18±0.81 21.22±1.05 31.13±0.71 37.30±0.93 46.84±1.33
OnPro + Ours (Ens.) 74.49±2.14 78.64±1.42 34.76±1.12 41.89±0.82 50.01±0.85 21.81±1.02 32.00±0.72 38.18±1.02 47.93±1.26
Table 9: Comparison of the final average accuracy achieved through independent inference and the use of the ensemble method on four benchmark datasets with difference memory buffer size $M$. All values are averages of 10 runs.
Dataset CIFAR10 CIFAR100 Tiny-ImageNet ImageNet-100
Memory Size $M$ 500 1000 1000 2000 5000 2000 5000 10000 5000
ER 33.16±3.50 20.94±6.79 32.65±1.78 22.20±2.26 13.29±1.98 58.38±1.69 46.87±1.60 40.77±2.45 23.38±2.10
ER + Ours 30.22±3.75 19.85±2.55 43.28±1.67 29.35±1.50 16.88±1.99 69.56±1.54 53.13±0.85 42.63±0.80 28.48±1.50
DER++ 24.21±2.75 18.42±1.84 34.49±4.39 25.55±3.26 20.01±2.88 62.03±2.83 51.57±4.60 49.51±3.04 28.77±4.10
DER++ + Ours 25.08±2.88 18.47±3.12 42.76±1.31 31.13±2.41 18.45±2.89 72.59±1.29 57.71±1.80 50.31±2.34 27.22±2.17
ER-ACE 12.72±3.56 10.66±2.48 12.67±1.62 9.11±0.78 5.92±1.09 19.12±0.63 17.14±0.67 15.59±1.24 14.11±1.19
ER-ACE + Ours 22.86±2.23 16.07±2.38 35.85±1.12 25.84±1.96 14.21±0.85 24.10±2.00 19.14±1.91 15.14±1.60 26.02±2.33
OCM 13.68±4.25 11.63±2.62 14.99±1.55 9.16±1.75 3.76±1.16 26.12±1.63 19.74±1.30 15.92±1.47 3.25±0.90
OCM + Ours 11.59±2.24 9.18±2.03 16.69±2.36 10.07±1.37 3.99±0.78 26.16±1.90 19.99±1.96 15.56±1.06 8.91±1.18
GSA 25.45±2.86 16.42±3.59 33.97±2.55 22.74±1.83 12.31±2.35 27.23±2.01 23.61±2.26 20.58±2.09 24.53±1.59
GSA + Ours 28.47±3.08 19.00±2.08 42.41±1.44 31.09±1.86 16.77±1.87 64.86±1.19 48.23±1.28 35.79±1.46 32.77±2.07
OnPro 17.94±3.69 14.20±2.60 16.76±2.47 12.42±1.39 6.72±0.94 28.01±1.59 23.52±1.75 20.32±1.70 7.59±1.17
OnPro + Ours 19.89±4.01 14.62±2.75 28.93±2.19 20.23±1.03 10.55±1.89 28.21±1.58 20.86±1.13 16.17±0.63 9.90±1.93
Table 10: Forgetting Measure (%, lower is better) on four benchmark datasets with difference memory buffer size $M$, with and without our proposed CCL-DC scheme. The result of our method is given by the ensemble of two peer models. All values are averages of 10 runs.
Performance with NCM classifier.
@@
Besides t-SNE, we can evaluate the feature discrimination using the clustering methods. We remove the final FC classifier and use Nearest-Class-Mean (NCM) classifier with intermediate representations. Table 11 demonstrates that CCL-DC can greatly enhance the NCM accuracy, which evidences the capability of CCL-DC in improving feature discrimination.
GradCAM++ visualization.
Shortcut learning is another commonly observed issue that hinders the generalization capability of continual learners [44]. In Fig. 7, we use GradCAM++ on the training set of ImageNet-100 ($M=$5k) at the end of the training of ER and GSA. Although both ER and GSA make correct predictions, we observed that they focus on irrelevant objects, which indicates a tendency toward shortcut learning. Also, we can see that by integrating CCL-DC, the shortcut learning can be greatly alleviated.
@@ 10 Experiment Details
All the experiments in our work are conducted on NVIDIA A100 GPUs. Fig. 12 shows the training time for each method with and without CCL-DC on CIFAR-100 ($M=$2k).